OpenStudy (animalover02):

How much compound interest is earned on $3500 at 3.6% interest for 7 years, compounded quarterly? Give your answer in dollars and cents. Don’t forget to subtract P from A. Compound Interest

OpenStudy (animalover02):

@jupitersring01

OpenStudy (animalover02):

@welshfella

OpenStudy (welshfella):

the interest will be 3.6 / 4 = 0.9% quarterly and number of payments over 7 years = 7*4 = 28 A = 35001 + 0.009)^28

OpenStudy (welshfella):

Subtract 3500 form the value of A to get the interest paid.

OpenStudy (animalover02):

i have 10 sec. left please just give me the answer

OpenStudy (welshfella):

i cant give answers you can do it quickly using ur calc

OpenStudy (animalover02):

35001-3500=31501<----correct? final answer<----

OpenStudy (welshfella):

3500*1.009^28

OpenStudy (animalover02):

what is the 28 for?

OpenStudy (animalover02):

so times 28?

OpenStudy (welshfella):

thats the power to which 1.009 is taken use the ^ key on your calc

OpenStudy (animalover02):

not on my cac.

OpenStudy (welshfella):

No it means 1.009 multiplied by itself 28 times

OpenStudy (animalover02):

so 1.009 multiplied by 28 is my final answer?

OpenStudy (welshfella):

u need a scientific calc There's one on your PC or laptop

OpenStudy (animalover02):

so 1.009^28 will be my final answer?

OpenStudy (welshfella):

n0 A comes to 3500 * 1.009 * 1.009 * 1.009........ (28 times)

OpenStudy (animalover02):

ok ill see what i get

OpenStudy (welshfella):

then take P (3500) from A

OpenStudy (animalover02):

i cant do it im keep getting off trac so ill just get someone elses help

OpenStudy (welshfella):

havent you got a calculator on your PC?

OpenStudy (animalover02):

no

OpenStudy (welshfella):

just google calculator

OpenStudy (welshfella):

Hold on

OpenStudy (welshfella):

https://www.google.co.uk/webhp?sourceid=chrome-instant&ion=1&espv=2&ie=UTF-8#q=calculator%20online Now enter 3500 X 1.009 EXP 28

OpenStudy (welshfella):

Sorry i'll do thet again 3500 X 1.009 xy 28 = The xy button is next to EXP

OpenStudy (animalover02):

i dont understand so i do 3500times1.009times 28?

OpenStudy (animalover02):

please do for me i cant

OpenStudy (animalover02):

wait nvmd your not allowed

OpenStudy (welshfella):

No after the 1.009 you press the button next to EXP It looks like 

OpenStudy (welshfella):

can you see it?

OpenStudy (animalover02):

no

OpenStudy (welshfella):

Its on the bottom line - just to the right of the EXP button

OpenStudy (welshfella):



OpenStudy (animalover02):

im gonna just get it wrong i dont have time for ths and mom needs me so i need answer from you or i will guess

OpenStudy (animalover02):

bye

OpenStudy (welshfella):

bye